# Dark lens candidates from Gaia Data Release 3

### v1

@@ Appendix A Additional tables
GaiaDR3- Survey, filter Num. of points GaiaDR3- Survey, filter Num. of points
ULENS- ULENS-
003 Gaia, $G$ 27 018 Gaia, $G$ 24
Gaia, $G_{\rm BP}$ 27
Gaia, $G_{\rm RP}$ 27 OGLE, I 537
ASASSN, V 84
023 Gaia, $G$ 64 025 Gaia, $G$ 33
Gaia, $G_{\rm BP}$ 60 Gaia, $G_{\rm RP}$ 32
Gaia, $G_{\rm RP}$ 60 OGLE, I 99
OGLE, I 125
ASASSN, V 410
032 Gaia, $G$ 40 035 Gaia, $G$ 6
Gaia, $G_{\rm BP}$ 38 Gaia, $G_{\rm RP}$ 5
Gaia, $G_{\rm RP}$ 37 OGLE, I 371
ASASSN, V 307 MOA 3448
047 Gaia, $G$ 21 057 Gaia, $G$ 24
Gaia, $G_{\rm RP}$ 10
KMTNet SAAO , I 804 Gaia, $G_{\rm RP}$ 23
KMTNet CTIO, I 363
KMTNet SSO, I 720
069 Gaia, $G$ 36 073
Gaia, $G_{\rm RP}$ 30 Gaia, $G$ 22
KMTNet SAAO, I 731 Gaia, $G_{\rm RP}$ 13
KMTNet CTIO, I 831 OGLE, I 604
KMTNet SSO, I 731 MOA 6102
MOA 29814
078 079 Gaia, $G$ 24
Gaia, $G_{\rm RP}$ 22
Gaia, $G$ 40 OGLE, I 1255
Gaia, $G_{\rm BP}$ 39 KMTNet SAAO, I 2010
Gaia, $G_{\rm RP}$ 39 KMTNet CTIO, I 4314
OGLE, I 634 KMTNet SSO, I 2669
MOA 9285
088 Gaia, $G$ 33 089 Gaia, $G$ 25
OGLE, I 1678 OGLE, I 609
KMTNet SAAO, I 511 KMTNet SAAO, I 249
KMTNet CTIO, I 819 KMTNet CTIO, I 395
KMTNet SSO, I 730 KMTNet SSO, I 348
MOA 24908
097 Gaia, $G$ 54 103 Gaia, $G$ 66
Gaia, $G_{\rm RP}$ 52 Gaia, $G_{\rm BP}$ 59
OGLE, I 543 Gaia, $G_{\rm RP}$ 63
OGLE, I 143
118 Gaia, $G$ 22 127 Gaia, $G$ 25
Gaia, $G_{\rm RP}$ 20 OGLE, I 1918
OGLE, I 166 KMTNet SAAO, I 285
ASASSN, V 19 KMTNet CTIO, I 412
KMTNet SSO, I 382
142 Gaia, $G$ 29 143 Gaia, $G$ 60
Gaia, $G_{\rm RP}$ 25 OGLE, I 122
OGLE, I 274
155 Gaia, $G$ 47 196
Gaia, $G_{\rm RP}$ 44
OGLE, I 1347 Gaia, $G$ 60
KMTNet SAAO, I 216 Gaia, $G_{\rm RP}$ 56
KMTNet CTIO, I 379 OGLE, I 572
KMTNet SSO, I 324
MOA 3290
212 Gaia, $G$ 59 230
GSA, $G$ 68
OGLE, I 189 Gaia, $G$ 28
KMTNet SAAO, I 706 OGLE, I 1985
KMTNet CTIO, I 971
KMTNet SSO, I 867
259 270 Gaia, $G$ 24
Gaia, $G$ 58 OGLE, I 778
Gaia, $G_{\rm RP}$ 50 KMTNet SAAO, I 568
GSA, $G$ 164 KMTNet CTIO, I 859
OGLE, I 52 KMTNet SSO, I 778
275 284 Gaia, $G$ 42
Gaia, $G$ 40 OGLE, I 643
Gaia, $G_{\rm RP}$ 39 KMTNet SAAO, I 640
OGLE, I 130 KMTNet CTIO, I 992
KMTNet SSO, I 888
326 Gaia, $G$ 39 331 Gaia, $G$ 42
OGLE, I 1625 OGLE, I 843
KMTNet SAAO, I 538 KMTNet SAAO, I 549
KMTNet CTIO, I 827 KMTNet CTIO, I 549
KMTNet SSO, I 746 KMTNet SSO, I 768
343 Gaia, $G$ 20 353 Gaia, $G$ 18
Gaia, $G_{\rm RP}$ 9 Gaia, $G_{\rm RP}$ 11
OGLE, I 1685 OGLE, I 742
KMTNet SAAO, I 650 KMTNet SAAO, I 625
KMTNet CTIO, I 908 KMTNet SSO, I 800
KMTNet SSO, I 835 MOA 3066
MOA 15768
359 Gaia, $G$ 21 363 Gaia, $G$ 23
OGLE, I 486 OGLE, I 457
KMTNet SAAO, I 212 KMTNet SAAO, I 758
KMTNet CTIO, I 382 KMTNet CTIO, I 1026
KMTNet SSO, I 343 KMTNet SSO, I 928

### v2

@@ Appendix A Additional tables
{ThreePartTable}{TableNotes}
Notes: The columns are event name, mass function MF used as a prior for lens mass, the mass of the lens $M_{\mathrm{L}}$ in solar masses, distance to the lens $D_{\mathrm{L}}$ in kpc, the brightness of an MS star with $M_{\mathrm{L}}$ mass at a distance $D_{\mathrm{L}}$, the brightness of the blend $G_{\mathrm{blend}}$ obtained from the microlensing model, Einstein radius $\theta_{\mathrm{E}}$, maximal astrometric displacement $\delta_{\mathrm{max}}$, proper motion of the lens $\mu_{\mathrm{L}}$, transverse velocity $v_{\mathrm{t}}$, and probability for a dark lens. In the MF column, K2001 refers to Kroupa (2001), and M2021 to Mróz et al. (2021).
{ThreePartTable}{TableNotes}
Notes: The table contains the event name, its equatorial coordinates (RA, dec), proper motion in Right Ascension $\mu_{\mathrm{RA}}$ and declination $\mu_{\mathrm{dec}}$ from Gaia DR3 and their correlation corr($\mu_{\mathrm{RA}}$, $\mu_{\mathrm{dec}}$), as well as the Renormalized Unit Weight Error (RUWE) (if available), and angular stellar radius $\theta_{*}$ found for the event solution. Column ”Flag” marks if the event’s colour is within the applicable range for the relation form Adams et al. (2018). Column ”Type” marks the assumed source star type for further analysis.
Table 3: Source star properties. The description of the columns is at the end of the table.
GaiaDR3- RA dec $\mu_{\mathrm{RA}}$ $\mu_{\mathrm{Dec}}$ corr($\mu_{\mathrm{RA}}$, $\mu_{\mathrm{dec}}$) RUWE $\theta_{*}$ Flag Type
ULENS- [deg] [deg] [$\mathrm{mas}\,\mathrm{yr}^{-1}$] [$\mathrm{mas}\,\mathrm{yr}^{-1}$] [$\mu\mathrm{mas}$]
025-G1- 260.8781 -27.3788 $-4.61\pm 0.06$ $-5.56\pm 0.04$ $0.21$ $1.08$ $5.58\pm 1.22$ true giant
025-G2- 260.8781 -27.3788 $-4.61\pm 0.06$ $-5.56\pm 0.04$ $0.21$ $1.08$ $6.22\pm 1.44$ true giant
035-G+ 267.9705 -34.9749 $-4.37\pm 0.11$ $-3.95\pm 0.07$ $0.23$ $1.83$ $6.49\pm 1.12$ true giant
035-G- 267.9705 -34.9749 $-4.37\pm 0.11$ $-3.95\pm 0.07$ $0.23$ $1.83$ $6.46\pm 1.11$ true giant
069-G+ 269.0712 -32.4500 $-2.94\pm 0.14$ $-3.11\pm 0.10$ $0.41$ $1.33$ $6.19\pm 1.03$ true giant
069-G- 269.0712 -32.4500 $-2.94\pm 0.14$ $-3.11\pm 0.10$ $0.41$ $1.33$ $6.34\pm 1.05$ true giant
073-G+ 272.4690 -30.0406 $1.18\pm 0.11$ $-7.59\pm 0.08$ $0.30$ $1.05$ $3.47\pm 0.83$ true giant
073-G- 272.4690 -30.0406 $1.18\pm 0.11$ $-7.59\pm 0.08$ $0.30$ $1.05$ $3.67\pm 0.78$ true giant
088-G- 268.1614 -32.4896 $-1.49\pm 0.09$ $-7.48\pm 0.07$ $0.17$ $0.92$ $5.39\pm 0.90$ true giant
089-G+ 270.4467 -26.5331 $1.67\pm 0.12$ $-8.26\pm 0.08$ $0.09$ $1.00$ $1.96\pm 3.34$ true dwarf
089-G- 270.4467 -26.5331 $1.67\pm 0.12$ $-8.26\pm 0.08$ $0.09$ $1.00$ $1.80\pm 3.07$ true dwarf
097-G+ 267.5594 -23.6717 $-4.15\pm 0.11$ $-4.53\pm 0.07$ $0.25$ $0.98$ $0.95\pm 1.75$ true dwarf
097-G- 267.5594 -23.6717 $-4.15\pm 0.11$ $-4.53\pm 0.07$ $0.25$ $0.98$ $0.94\pm 1.73$ true dwarf
103-G+ 249.8323 -40.8217 $-5.04\pm 0.13$ $-5.43\pm 0.09$ $-0.38$ $1.00$
103-G- 249.8323 -40.8217 $-5.04\pm 0.13$ $-5.43\pm 0.09$ $-0.38$ $1.00$
118-G+ 99.4167 13.9551 $1.34\pm 0.23$ $-0.11\pm 0.19$ $-0.50$ $2.07$
118-G- 99.4167 13.9551 $1.34\pm 0.23$ $-0.11\pm 0.19$ $-0.50$ $2.07$
142-G1+ 263.5395 -24.8825 $-4.15\pm 0.11$ $-4.53\pm 0.07$ $0.25$ $1.09$ $2.72\pm 4.69$ true dwarf
142-G2+ 263.5395 -24.8825 $-4.15\pm 0.11$ $-4.53\pm 0.07$ $0.25$ $1.09$ $2.72\pm 4.69$ true dwarf
142-G2- 263.5395 -24.8825 $-4.15\pm 0.11$ $-4.53\pm 0.07$ $0.25$ $1.09$ $3.14\pm 5.17$ true dwarf
155-G+ 270.3789 -32.6859 $-0.69\pm 0.16$ $-7.97\pm 0.11$ $0.35$ $1.02$ $2.76\pm 0.55$ true giant
155-G1- 270.3789 -32.6859 $-0.69\pm 0.16$ $-7.97\pm 0.11$ $0.35$ $1.02$ $2.89\pm 0.54$ true giant
155-G2- 270.3789 -32.6859 $-0.69\pm 0.16$ $-7.97\pm 0.11$ $0.35$ $1.02$ $2.46\pm 0.49$ true giant
212-GSA+ 265.6345 -34.1037 $-1.91\pm 0.17$ $-4.99\pm 0.09$ $0.22$ $1.07$ $2.31\pm 0.39$ true dwarf
212-GSA- 265.6345 -34.1037 $-1.91\pm 0.17$ $-4.99\pm 0.09$ $0.22$ $1.07$ $2.31\pm 0.39$ true dwarf
259-GSA+ 116.0932 -28.4434 $-1.54\pm 0.20$ $2.72\pm 0.23$ $0.17$ $2.00$
259-GSA- 116.0932 -28.4434 $-1.54\pm 0.20$ $2.72\pm 0.23$ $0.17$ $2.00$
270-G+ 264.4791 -28.5589 $-3.18\pm 0.44$ $-1.22\pm 0.31$ $0.30$ $1.03$ $2.62\pm 4.43$ true dwarf
270-G- 264.4791 -28.5589 $-3.18\pm 0.44$ $-1.22\pm 0.31$ $0.30$ $1.03$ $2.62\pm 4.45$ true dwarf
331-G+ 264.9230 -27.6781 $-3.99\pm 0.82$ $-3.64\pm 0.48$ $0.42$ $1.08$ $1.12\pm 1.89$ true dwarf
331-G- 264.9230 -27.6781 $-3.99\pm 0.82$ $-3.64\pm 0.48$ $0.42$ $1.08$ $1.12\pm 1.89$ true dwarf
343-G+ 267.8664 -33.1352 $0.87\pm 0.14$ true dwarf
343-G- 267.8664 -33.1352 $0.87\pm 0.14$ true dwarf
353-G+ 267.6884 -34.2448 $0.54\pm 0.09$ true dwarf
353-G- 267.6884 -34.2448 $0.54\pm 0.09$ true dwarf
363-G+ 266.9194 -25.0204 $1.09\pm 1.93$ false dwarf
363-G- 266.9194 -25.0204 $1.09\pm 1.94$ false dwarf
[…]